# Generative Diffusion Model Driven Massive Random Access in Massive MIMO Systems

### v1

@@ V-B1 Calculation of Prior Score Function
where the inner expectation is taken over the joint distribution $p(\bm{h}(t),\bm{h}(0))$ with $\bm{h}(0)$ sampled from the dataset distribution $p_{0}(\bm{h})$, and the time $t$ is uniformly sampled from (0, 1]01(0,\,1]( 0 , 1 ] with $T\!=\!1$, while $\lambda(t)\!\propto\!1/\mathbb{E}[\|\nabla_{\bm{h}(t)}\log p(\bm{h}(t)|\bm{h}(0))\|_{2}^{2}]$ is a positive weight function, and $\bm{\theta}$ collects the neural network parameters.
@@ V-B3 Channel Sampling Process
[…]
@@
×$\hskip 93.89409pt\times\left(\bm{U}^{\rm T}\bm{Y}-\bm{\Sigma}\bm{V}^{\rm T}\bm{H}_{a}^{i}\right),$ (24b)
@@ V-B4 Overall CE Algorithm
0:  ${\bm{S}_{a}}$, ${\bm{Y}}$, $\sigma_{n}^{2}$, pre-trained denoising network $\bm{s}_{\bm{\theta}^{*}}$, step number $N$, noise scheduling function $\sigma(t)$, and hyper-parameter $\lambda_{h}$;
@@
6:     ${\bm{H}}_{a}^{i-1}=\bm{H}_{a}^{i}+\left(\sigma_{i}^{2}-\sigma_{i-1}^{2}\right)\nabla_{\bm{H}_{a}^{i}}\log p(\bm{H}_{a}^{i}|\bm{Y})+\sqrt{\sigma_{i}^{2}-\sigma_{i-1}^{2}}\bm{Z}$;
@@ V-C Data Detection
×$\times\left(\bm{U}_{h}^{T}\mathring{\bm{Y}}_{d}-\bm{\Sigma}_{h}\bm{V}_{h}^{\rm T}\mathring{\bm{X}}_{a}^{i}\right),$ (30)
@@ V-D Overall Algorithm
    $i^{*}=\arg\min\limits_{\forall i}\bigg{|}\sigma_{i}-\sum_{l=1}^{K_{a}M}\left[\bm{R}_{\Delta}\right]_{l,l}\bigg{|}.$ (32)
@@
0:  ${\bm{S}_{a}}$, ${\bm{Y}}$, $\sigma_{n}^{2}$, pre-trained denoising network $\bm{s}_{\bm{\theta}^{*}}$, SDE noise scheduling functions $\sigma(t)$ and $\tau(t)$, step numbers $N_{H}$ and $N_{D}$, hyper-parameters $\lambda_{h}$ and $\lambda_{x}$;
@@ V-E Computational Complexity Analysis
The complexity analysis of Algorithm 2 is detailed in Table I. The channel is initialized using the LMMSE method before the iterative process begins. For the channel PC-sampler, $i^{\star}$ iterations are necessary. Each PC iteration consists of one predictor step and $Q$ corrector steps, resulting in a total of $i^{\star}(1\!+\!Q)$ computations of posterior scores, as described in Lines 4-5 of Algorithm 2. Furthermore, data is updated after every $N_{\text{update}}$ channel iterations, as outlined in Lines 6-14.
@@
For typical simulation parameters, e.g., $N_{\text{update}}\!=\!50$, $\big{\lfloor}\frac{i^{\star}}{N_{\text{update}}}\big{\rfloor}\!=\!5\!\sim\!15$, $Q\!=\!3$, $N_{X}\!=\!1500$ and $N_{H}\!=\!1500$, the algorithm’s overall complexity is mainly determined by the score function computations in Lines 4-5 and Lines 10-11. In subsequent simulations, the computational delay of the proposed algorithm will be compared with those of other baseline algorithms to verify its complexity efficacy.
@@ VI-A Experimental Setups
We consider a typical uplink random access scenario with a total of $K\!=\!128$ potential users, where each user’s activity follows a Bernoulli distribution with an active probability of $0.1$. The BS employs a UPA with $M_{x}\!=\!M_{y}\!=\!8$ antennas on each dimension, while each user operates with a single antenna. Each element of the pilot sequence $\tilde{\bm{p}}_{k}$, $\forall k$, and the data symbols $\bm{x}_{k}$ are randomly sampled from 4QAM constellations. The maximum pilot sequence length is $L_{p}^{\text{max}}\!=\!28$. We consider a 3GPP-compliant channel model [33] generated using Quadriga [41], simulating the 3GPP 38.901 UMa NLOS scenario at a carrier frequency of 3.5 GHz. The minimum pilot sequence length is $L_{p}^{\text{min}}\!=\!8$ for Rayleigh channel, and $L_{p}^{\min}\!=\!10$ for 3GPP channel, since 3GPP channel is more complicated and it needs more pilots to estimate AUS.
@@ VI-A1 VPL-AUDNet Setup
A total of 64,000 channel samples are generated using Quadriga. For each training batch, the pilot lengths for the UEs are randomly selected from $[L_{p}^{\text{min}},\,L_{p}^{\text{max}}]$. In addition to the input preprocessing layer and output layer, the VPL-AUDNet contains 5 heterogeneous Transformer encoder layers, each featuring a hidden state sequence length of $d\!=\!512$ and 8 attention heads. Training is conducted over 200 epochs, with the learning rate warming up to $10^{-4}$ during the initial 15 epochs and then decaying to 0.1 of this original value every 75 epochs. The batch size is set to 128.
@@ VI-A2 Channel Prior Score Network Setup
The NSCN++ architecture [32] is employed to model the prior score function $\bm{s}_{\bm{\theta}}(\bm{h}(t),t)$. A 3GPP training dataset consisting of 160,000 channel samples is utilized. The time step $t$ is uniformly sampled from the interval [0, 1]01[0,\,1][ 0 , 1 ], and the perturbation noise variance is determined using $\sigma(t)\!=\!\sigma_{\min}\left(\frac{\sigma_{\max}}{\sigma_{\min}}\right)^{t}$, with $\sigma_{\text{max}}\!=\!30$ and $\sigma_{\text{min}}\!=\!0.01$. The network is trained with a learning rate of $10^{-5}$ and a batch size of 64, over 20 epochs.
@@ VI-B Performance of AUD
Fig. 6(a) illustrates the AUD performance of various schemes as the functions of the transmission pilot length, $L_{p}^{\text{test}}$, under i.i.d. Rayleigh channels, where the channel between each potential UE and the BS is modeled as $\bm{h}_{k}\!\sim\!\mathcal{N}_{c}(\bm{h}_{k};\bm{0},\bm{I}_{M}),\forall k$. The simulation is conducted under the received $\text{SNR}\!=\!\frac{\mathbb{E}\left\{\|\bm{S}\bm{H}\|^{2}\right\}}{\mathbb{E}\left\{\|\bm{W}\|^{2}\right\}}\!=\!10$ dB. The results indicate that that the VPL-AUDNet-based schemes outperform traditional CS algorithms based on AMP and OMP, in terms of AEP. Furthermore, the VPL-AUDNet models trained with fixed pilot lengths and without the PLAM module, e.g., VPL-AUDNet (w/o PLAM), $L_{p}^{\text{train}}\!=\!15$, exhibit degraded performance when tested on pilot lengths different from the training pilot length. Conversely, VPL-AUDNet (w/o PLAM), $L_{p}^{\text{train}}\!\in\![10,\,28]$ exhibit better generalization performance. Moreover, the proposed VPL-AUDNet (w/ PLAM), $L_{\text{train}}\!\in\![8,\,24]$, outperforms VPL-AUDNet (w/o PLAM), $L_{\text{train}}\!\in\![8,\,24]$, demonstrating the effectiveness of the PLAM. Additionally, under i.i.d. Rayleigh channels, the proposed VPL-AUDNet scheme achieves an AEP performance close to $10^{-2}$ even when the pilot number is less than the average number of UEs ($L_{p}^{\text{test}}\!\leq\!12$), whereas the AMP and OMP algorithms exhibit a AEP of $10^{-1}$. These results demonstrate that the proposed scheme achieves optimal overall performance across various pilot lengths within a single model, thereby confirming its effectiveness.
@@
Fig. 7 depicts the curves for the probability of false alarm ($P_{\text{fa}}$) and the probability of missed detection ($P_{\text{md}}$) for various schemes under 3GPP channels. Simulations are conducted with a received SNR of 10 dB and a test pilot length of $L_{p}^{\text{test}}\!=\!25$. All VPL-AUDNet models are trained with an antenna number of $M\!=\!64$. The results indicate that the proposed VPL-AUDNet (w/ PLAM, w/ SCM), $L_{p}^{\text{train}}\!=\![10,\,28]$, achieves an optimal trade-off between $P_{\text{fa}}$ and $P_{\text{md}}$. Moreover, testing the model trained with $M\!=\!64$ antennas on the channels with $M\!=\!144$ antennas results in further improvements in $P_{\text{fa}}\text{-}P_{\text{md}}$ performance, demonstrating the scheme’s robust generalization capability across varying antenna configurations. In contrast, the performance gains for OMP related to an increased number of antennas are less significant.
@@ VI-C Performance of JCEDD
In the simulations, unless otherwise specified, the BS is equipped with $M\!=\!64$ antennas, and a total of $K_{a}\!=\!12$ active UEs access the system simultaneously.
@@
Figure 8: NMSE and BER performance with respect to the received SNR for different CE and DD algorithms.
@@
[…]
@@
Fig. 12 compares the computational latency of various methods for processing a batch size of 64 in CE and DD tasks, utilizing a NVIDIA GeForce GTX 4090 GPU and an AMD EPYC 7J13 CPU. Due to its asynchronous update mechanism and LMMSE initialization, Our Iter SDE CE & DD algorithm demonstrates significantly lower computational latency compared to Iter Langevin CE & DD algorithm. Furthermore, the latency of Our Iter SDE CE & DD is comparable to that of Iter LMMSE-CE & OAMP-DD, while the previous simulations have shown it outperforming Iter LMMSE-CE & OAMP-DD in both CE and DD. Although the latency of iterative CE and DD algorithms is higher than that of two-stage algorithms, such as Pilot LS-CE & ZF-DD and Pilot LS-CE & OAMP-DD, existing researches indicate that diffusion-based algorithms can reduce the number of steps from thousands to tens, or even to as few as 1 to 2 steps, through acceleration methods such as higher-order Langevin algorithms [26] and consistency models [42]. Therefore, accelerating diffusion algorithms still holds significant potential for addressing CE and DD problems.

### v2

@@ V-B1 Calculation of Prior Score Function
where the inner expectation is taken over the joint distribution $p(\bm{h}(t),\bm{h}(0))$ with $\bm{h}(0)$ sampled from the dataset distribution $p_{0}(\bm{h})$, and the time $t$ is uniformly sampled from (0, 1](0,\,1] with T=1T\!=\!1, while $\lambda(t)\!\propto\!1/\mathbb{E}[\|\nabla_{\bm{h}(t)}\log p(\bm{h}(t)|\bm{h}(0))\|_{2}^{2}]$ is a positive weight function, and $\bm{\theta}$ collects the neural network parameters.
@@
One special case where the channel prior score function does not need to be learned is when the channel distribution has a known closed form, for example, $\bm{h}(0)\sim\mathcal{N}\left(\bm{h}(0);\bm{\mu}_{h},\bm{R}_{hh}\right)$. Then, the prior score function of the channel in this case can be derived by Tweedie’s formula [45]
$\nabla_{\bm{h}(t)}\log p(\bm{h}(t))$ $=\frac{\mathbb{E}\left\{\bm{h}(0)|\bm{h}(t)\right\}-\bm{h}(t)}{\sigma^{2}(t)}$ (19)
$=-\left(\bm{R}_{hh}+\sigma(t)^{2}\bm{I}\right)^{-1}\left(\bm{h}(t)-\bm{\mu}_{h}\right),$
where the posterior expectation can be derived according to Bayes linear Gaussian model, which is given by $\mathbb{E}\left\{\bm{h}(0)|\bm{h}(t)\right\}=\bm{\mu}_{h}+\bm{R}_{hh}\left(\bm{R}_{hh}+\sigma^{2}(t)\bm{I}\right)^{-1}\left(\bm{h}(t)-\bm{\mu}_{h}\right)$.
@@ V-B3 Channel Sampling Process
Utilizing the derived prior score function and likelihood score function, we numerically solve the reverse SDE (V-B). Specifically, by dividing the interval $t\!\in\!(0,\,1]$ into $N$ discrete steps, indexed as $i\!=\!0,1,\dots,N-1$, we obtain the following iterative rule for the reverse diffusion sampler (which is also called predictor):
@@
$\hskip 93.89409pt\times\left(\bm{U}^{\rm T}\bm{Y}-\bm{\Sigma}\bm{V}^{\rm T}\bm{H}_{a}^{i}\right),$ (25b)
@@ V-B4 Overall CE Algorithm
[…]
@@
6:  ${\bm{H}}_{a}^{i-1}=\bm{H}_{a}^{i}+\left(\sigma_{i}^{2}-\sigma_{i-1}^{2}\right)\nabla_{\bm{H}_{a}^{i}}\log p(\bm{H}_{a}^{i}|\bm{Y})+\sqrt{\sigma_{i}^{2}-\sigma_{i-1}^{2}}\bm{Z}$; % Predictor: line 6
@@ V-C Data Detection
$\times\left(\bm{U}_{h}^{T}\mathring{\bm{Y}}_{d}-\bm{\Sigma}_{h}\bm{V}_{h}^{\rm T}\mathring{\bm{X}}_{a}^{i}\right),$ (31)
@@ V-D Overall Algorithm
    $i^{*}=\arg\min\limits_{\forall i}\bigg|\sigma_{i}-\frac{1}{K_{a}M}\sum_{l=1}^{K_{a}M}\left[\bm{R}_{\Delta}\right]_{l,l}\bigg|.$ (33)
@@
[…]
@@ V-E Computational Complexity Analysis
The complexity analysis of Algorithm 2 is detailed in Tab. 1. The channel is initialized using the LMMSE method before the iterative process begins. For the channel PC-sampler, $i^{\star}$ iterations are necessary. Each PC iteration consists of one predictor step and $Q$ corrector steps, resulting in a total of $i^{\star}(1\!+\!Q)$ computations of posterior scores, as described in Lines 4-5 of Algorithm 2. Furthermore, data is updated after every $N_{\text{update}}$ channel iterations, as outlined in Lines 6-14.
@@
For typical simulation parameters, e.g., $N_{\text{update}}\!=\!50$, $\big\lfloor\frac{i^{\star}}{N_{\text{update}}}\big\rfloor\!=\!5\!\sim\!15$, Q=3Q\!=\!3, $N_{X}\!=\!1500$ and $N_{H}\!=\!1500$, the overall complexity of Algorithm 2 is primarily determined by the score function computations in lines 4-5 and 10-11. Specifically, the neural network NCSN++ [38], employed to calculate the channel’s prior score function, markedly impacts the inference complexity. For complex channel distributions, utilizing a moderately complex network can balance performance with computational demands, providing an avenue for algorithm optimization. Additionally, given that the proposed approach includes multiple PC-sampler iterations, the use of higher-order SDE or ordinary differential equation (ODE) solvers [47, 48] is essential for further reducing complexity in engineering applications.
@@ VI-A Experimental Setups
We consider a typical uplink random access scenario with a total of K=128K\!=\!128 potential users [^1] Future networks are expected to support millions of devices per square kilometer. However, such high user density can only be achieved within specific time-frequency resources. For anticipated 6G massive communications, it is generally necessary to divide limited resources into different time-frequency resource groups. For example, within the coverage area of a single macro base station (approximately $S=\pi r^{2}=\pi\times 0.5^{2}=0.7854\,\text{km}^{2}$), 128 users can be grouped to share the same resource block, such as 12 subcarriers and 14 OFDM symbols. Considering a system with 1 frame = 10 ms = 140 OFDM symbols and 512 resource blocks (providing approximately 92.16 MHz bandwidth for 15 kHz subcarrier spacing), the system can theoretically support up to $512\times 10\times 128=655,360$ UEs, corresponding to a user density of $\frac{655,360\,\text{UEs/sector}\times 3\,\text{sectors}}{0.7854\,\text{km}^{2}}\approx 2.5\times 10^{6}\,\text{UEs/km}^{2}$ for the $100\,\text{MHz}\times 10\,\text{ms}$ time-frequency resources, effectively meeting the massive communication requirement. , where each user’s activity follows a Bernoulli distribution with an active probability of $0.1$. The BS employs a UPA with $M_{x}\!=\!M_{y}\!=\!8$ antennas on each dimension, while each user operates with a single antenna. Each element of the pilot sequence $\tilde{\bm{p}}_{k}$, $\forall k$, and the data symbols $\bm{x}_{k}$ are randomly sampled from 4QAM constellations. The maximum pilot sequence length is $L_{p}^{\text{max}}\!=\!28$. We consider a 3GPP-compliant channel model [39] generated using Quadriga [49], simulating the 3GPP 38.901 UMa NLOS scenario at a carrier frequency of 3.5 GHz. The minimum pilot sequence length is $L_{p}^{\text{min}}\!=\!8$ for Rayleigh channel, and $L_{p}^{\min}\!=\!10$ for 3GPP channel, since 3GPP channel is more complicated and it needs more pilots to estimate AUS.
@@ VI-A1 VPL-AUDNet Setup
A total of 64,000 channel samples are generated using Quadriga. For each training batch, the pilot lengths for the UEs are randomly selected from $L_{p}^{\text{min}},\,L_{p}^{\text{max}}]$. In addition to the input preprocessing layer and output layer, the VPL-AUDNet contains 5 heterogeneous Transformer encoder layers, each featuring a hidden state sequence length of d=512d\!=\!512 and 8 attention heads. Training is conducted over 200 epochs, with the learning rate warming up to $10^{-4}$ during the initial 15 epochs and then decaying to 0.1 of this original value every 75 epochs. The batch size is set to 128.
@@ VI-A2 Channel Prior Score Network Setup
The NSCN++ architecture [38] is employed to model the prior score function $\bm{s}_{\bm{\theta}}(\bm{h}(t),t)$. A 3GPP training dataset consisting of 160,000 channel samples is utilized. The time step $t$ is uniformly sampled from the interval [0, 1][0,\,1], and the perturbation noise variance is determined using $\sigma(t)\!=\!\sigma_{\min}\left(\frac{\sigma_{\max}}{\sigma_{\min}}\right)^{t}$, with $\sigma_{\text{max}}\!=\!30$ and $\sigma_{\text{min}}\!=\!0.01$. The network is trained with a learning rate of $10^{-5}$ and a batch size of 64, over 20 epochs.
@@ VI-B Performance of AUD
Fig. 6(a) illustrates the AUD performance of various schemes as the functions of the transmission pilot length, $L_{p}^{\text{test}}$, under i.i.d. Rayleigh channels, where the channel between each potential UE and the BS is modeled as $\bm{h}_{k}\!\sim\!\mathcal{N}_{c}(\bm{h}_{k};\bm{0},\bm{I}_{M}),\forall k$. The simulation is conducted under the received $\text{SNR}\!=\!\frac{\mathbb{E}\left\{\|\bm{S}\bm{H}\|^{2}\right\}}{\mathbb{E}\left\{\|\bm{W}\|^{2}\right\}}\!=\!10$ dB. The results indicate that that the VPL-AUDNet-based schemes outperform traditional CS algorithms based on AMP and OMP, in terms of AEP. Furthermore, the VPL-AUDNet models trained with fixed pilot lengths and without the PLAM module, e.g., VPL-AUDNet (w/o PLAM), $L_{p}^{\text{train}}\!=\!15$, exhibit degraded performance when tested on pilot lengths different from the training pilot length [^2] For models trained with a pilot length of $L_{p}^{\text{train}}$, if the test length is shorter than the training length, the received signals are padded with zeros to align with $L_{p}^{\text{train}}$ before being input into the VPL-AUDNet network. Conversely, if the test length exceeds the training length, the received signals are truncated to $L_{p}^{\text{train}}$ before network input.. Conversely, VPL-AUDNet (w/o PLAM), $L_{p}^{\text{train}}\!\in\![10,\,28]$ exhibit better generalization performance. Moreover, the proposed VPL-AUDNet (w/ PLAM), $L_{\text{train}}\!\in\![8,\,24]$, outperforms VPL-AUDNet (w/o PLAM), $L_{\text{train}}\!\in\![8,\,24]$, demonstrating the effectiveness of the PLAM. Additionally, under i.i.d. Rayleigh channels, the proposed VPL-AUDNet scheme achieves an AEP performance close to $10^{-2}$ even when the pilot number is less than the average number of UEs ($L_{p}^{\text{test}}\!\leq\!12$), whereas the AMP and OMP algorithms exhibit a AEP of $10^{-1}$. These results demonstrate that the proposed scheme achieves optimal overall performance across various pilot lengths within a single model, thereby confirming its effectiveness.
@@
Fig. 7 depicts the curves for the probability of false alarm ($P_{\text{fa}}$) and the probability of missed detection ($P_{\text{md}}$) for various schemes under 3GPP channels. Simulations are conducted with a received SNR of 10 dB and a test pilot length of $L_{p}^{\text{test}}\!=\!25$. All VPL-AUDNet models are trained with an antenna number of M=64M\!=\!64. The results indicate that the proposed VPL-AUDNet (w/ PLAM, w/ SCM), $L_{p}^{\text{train}}\!=\![10,\,28]$, achieves an optimal trade-off between $P_{\text{fa}}$ and $P_{\text{md}}$. Moreover, testing the model trained with M=64M\!=\!64 antennas on the channels with M=144M\!=\!144 antennas results in further improvements in $P_{\text{fa}}\text{-}P_{\text{md}}$ performance, demonstrating the scheme’s robust generalization capability across varying antenna configurations. In contrast, the performance gains for OMP related to an increased number of antennas are less significant.
Figure 8: AEP performance with respect to the active probability of different AUD schemes under the 3GPP channel.
Fig. 8 compares the AEP performance curves of different schemes as a function of active probability of UEs $p_{a}$ under the 3GPP channel. The simulation assumes a received SNR of 10 dB, a total potential number of UEs of 128, a user activation probability of $[0.06,0.2]$, and a test pilot length of $L_{p}^{\text{test}}=25$. The simulation results indicate that although AMP achieves marginally better AEP performance when the active probability is low, its performance deteriorates more rapidly as the active probability increases. In contrast, our scheme based on VPL-AUDNet exhibits a more gradual performance degradation, indicating better generalization capabilities in adapting to variations in activity probability.
Figure 9: Computation latency of AUD schemes under different pilot lengths.
Fig. 9 illustrates the computational latency required to process a batch of 128 samples across different schemes with varying pilot lengths. The results show that the computational latency of the VPL-AUDNet scheme remains approximately constant regardless of the pilot length due to a specialized receiver structure accommodating variable pilot length inputs. In contrast, as the pilot length increases, compressed sensing algorithms based on AMP or OMP experience increased computational complexity because of the corresponding growth in the dimension of the sensing matrix. The overall computational delay of the proposed VPL-AUDNet scheme is lower than that of the AMP scheme but higher than that of the OMP scheme. Nevertheless, the AUD performance of the VPL-AUDNet scheme is generally superior to that of OMP, highlighting the advantages of a neural network-based approach.
@@ VI-C Performance of JCEDD
In this subsection’s simulations, unless otherwise specified, the BS is equipped with M=64M\!=\!64 antennas, and a total of $K_{a}\!=\!12$ active UEs access the system simultaneously. Moreover, apart from Fig. 11, which utilizes a Rayleigh channel model, all other simulations are conducted using the 3GPP channel model.
@@
Figure 10: NMSE and BER performance with respect to the received SNR for different CE and DD algorithms under 3GPP channels.
@@
Figure 11: NMSE and BER performance with respect to the received SNR for different CE and DD algorithms under the MIMO Rayleigh channel.
In Fig. 11, we conducted simulations using the MIMO Rayleigh channel model, where the channel prior is assumed to follow $\mathcal{N}(\bm{0},\bm{I})$. Instead of utilizing a neural network-learned score, we substituted the derived prior score (19) into the process of Algorithm 2. The results indicate that Our Iter SDE CE & DD scheme outperforms traditional two-stage algorithms such as Pilot LS-CE & ZF-DD and Pilot LS-CE & OAMP-DD. However, it is less effective compared to the Iter LMMSE-CE & OAMP-DD algorithms. This is because, in a simple Rayleigh channel distribution, the mean and covariance matrix—representing first and second-order statistical information—are adequate to capture the channel’s distribution characteristics. Consequently, the LMMSE-CE and OAMP-DD achieve near-optimal performance, while the prior score offers limited additional information. Additionally, numerical errors in solving the SDE process can affect the scheme’s final performance. This suggests that the proposed scheme excels with more complex channel distributions, whereas existing methods already perform well in simple Rayleigh channels.
@@
Fig. 14 compares the performance of CE and DD of Our Iter SDE CE & DD scheme as the function of the maximum algorithm iteration numbers $N_{H}$ and $N_{X}$. In the simulation, each UE’s pilot length is set to $L_{p}\!=\!15$ and the data length is $L_{d}\!=\!50$, with the received SNR fixed at 10 dB. For the predefined noise diffusion functions $\sigma(t)$ and $\tau(t)$, the interval $t\!\in\![0,\,1]$ can be divided into segments with different values of $N_{X}$ and $N_{H}$ to control the noise variance gap between adjacent steps, thereby balancing generation speed and performance. As shown in Fig. 14(a), as the channel sampling steps $N_{H}$ increase, the NMSE performance improves gradually due to a more elaborate denoising process, resulting in more accurate SDE solutions. When $N_{H}\!=\!1000$, the NMSE of Our Iter SDE CE & DD approximates that of SDE-CE & Perfect-Data. Moreover, under the same $N_{H}$, CE performance remains similar for different $N_{D}$, possibly due to SDE-based CE not being sensitive to DD errors. Additionally, Fig. 14(b) shows that as the data sampling steps $N_{D}$ increase, the BER performance of DD gradually improves. However, the BER curves vary for different $N_{H}$, indicating that for DD problems, SDE-based DD methods are more sensitive to CE errors.
Figure 15: Convergence performance of (a) NMSE with varying pilot lengths $L_{p}$ and (b) BER under different SNRs, both as a function of the iteration index ii.
Fig. 15(a) illustrates the NMSE performance curve of Our Iter SDE CE & DD relative to the iteration index $i$ with varying pilot lengths $L_{p}$. The maximum number of iterations is set to $N_{H}=1500$ ($i\leq 1500$). The simulation results indicate that as the pilot length $L_{p}$ increases, the convergence iteration count for the proposed algorithm decreases from approximately 400 iterations at $L_{p}=12$ to about 100 iterations at $L_{p}=27$. Additionally, the final converged NMSE improves from 17.1-17.1 dB to 18.5-18.5 dB. Likewise, Fig. 15(b) demonstrates that with different SNRs, both the convergence speed and the final DD outcome of our algorithm improve as the SNR increases. Despite employing a maximum of 1500 iterations, the simulation results reveal that under favorable conditions, such as $L_{p}=27$ and $\text{SNR}=16$ dB, the proposed scheme requires only about 50 to 100 iterations to achieve satisfactory results.
Tab. 2: Computation latency of Algorithm 2 at different pilot lengths
Metrics Methods $L_{p}=12$ $L_{p}=15$ $L_{p}=18$ $L_{p}=21$ $L_{p}=24$ $L_{p}=27$
Latency [s] w/ LMMSE init 2.32 2.14 1.99 1.87 1.76 1.70
w/o LMMSE init 4.89 4.88 4.88 4.87 4.86 4.88
relative time cost 47.41% 43.86% 40.84% 38.42% 36.14% 34.79%
Tab. 3: Computation latency of Algorithm 2 at different SNRs
Metrics Methods $\text{SNR}=0$ dB $\text{SNR}=4$ dB $\text{SNR}=8$ dB $\text{SNR}=12$ dB $\text{SNR}=16$ dB
Latency [s] w/ LMMSE init 2.62 2.39 2.16 1.84 1.44
w/o LMMSE init 4.87 4.89 4.88 4.89 4.88
relative time cost 53.76% 48.82% 44.37% 37.61% 29.52%
In Tab. 2, we present the average computation latency for one single realization of Our Iter SDE CE & DD algorithm, both with and without LMMSE initialization, across various pilot lengths. Simulations were performed using an NVIDIA GeForce GTX 4090 GPU and an AMD EPYC 7J13 CPU. As the pilot length $L_{p}$ increases, the computational latency of the method with LMMSE initialization gradually decreases, whereas the latency without LMMSE initialization remains relatively constant. This occurs because a longer pilot results in more accurate initial channel estimation, which corresponds to a smaller diffusion noise-level in the iteration process, resulting in fewer required iteration steps. Similarly, Tab. 3 illustrates the differences in computational latency across various SNRs when LMMSE initialization is applied or not. Notably, our approach requires only 29%-53% of the original scheme’s complexity, underscoring the effectiveness of the LMMSE initialization strategy.
@@
Fig. 16 compares the computational latency of various methods for processing a batch size of 64 in CE and DD tasks. Due to its asynchronous update mechanism and LMMSE initialization, Our Iter SDE CE & DD algorithm demonstrates significantly lower computational latency compared to Iter Langevin CE & DD algorithm. Furthermore, the latency of Our Iter SDE CE & DD is comparable to that of Iter LMMSE-CE & OAMP-DD, while the previous simulations have shown it outperforming Iter LMMSE-CE & OAMP-DD in both CE and DD. Although the latency of iterative CE and DD algorithms is higher than that of two-stage algorithms, such as Pilot LS-CE & ZF-DD and Pilot LS-CE & OAMP-DD, existing researches indicate that diffusion-based algorithms can reduce the number of steps from thousands to tens, or even to as few as 1 to 2 steps, through acceleration methods such as higher-order algorithms [48] and consistency models [50]. Therefore, accelerating diffusion algorithms still holds significant potential for addressing CE and DD problems.